# Hypergraph adjusted plus-minus

### v1

@@ 1 Introduction
One of the most common approaches to measuring an individual’s effect in team sports is box-score plus-minus (PM). Originally used in hockey, a player’s PM is their team’s cumulative difference in score when that player is in the game. While intuitively appealing, PM is sensitive to substitution patterns. For instance, one very good player can artificially inflate the PM of their teammates, and quantifying this effect becomes challenging when certain combinations of players consistently share playing time. The most common approach to account for this is known as adjusted PM (APM). APM was developed for basketball [1] and hockey [2] as a method for comparing individual players while accounting for those that play together. We provide the mathematical details of APM in Section 3.
The intention of APM and its variants [3] is to isolate an individual player’s effect. However, this approach fails to capture any interaction effects between players. Experience, complimentary skill sets, and general synergy often result in a group of players whose combined effect surpasses the sum of their individual contributions. This phenomena is in part why substitution patterns are not uniformly random, and why managers and coaches play a crucial role in the success of a team. Likewise, when assessing and ranking an individual’s contribution to a team’s success, it is critical to account for the player’s ability to create synergies with the various combinations of their teammates. We refer to these various combinations (individuals, pairs, trios, etc.) as generalized lineups and propose a novel approach to quantify the contribution of all generalized lineups for a team simultaneously.
@@
Both of these approaches have obvious drawbacks, and neither is able to rank the generalized lineups while accounting for interaction effects. To fill the gap between these two extreme solutions, we introduce a new class of methods for ranking lineups and players that adjusts not only for a full lineup, but all lower-order combinations of players. Our methods are motivated by the hypergraph representation of a lineup, which we show is a generalization of the design matrix used in the ordinary APM method. The results are a unified approach for simultaneously ranking any generalized lineup including individuals, pairs, and full lineups. Furthermore, our methods lend themselves to interpretable visualizations that allow one to quickly identify positive and negative synergies among groups of players on a team.
@@ 2 Related Work
While methods for ranking lower-order combinations of players and full lineups simultaneously are lacking, there has been a great deal of research on measuring the efficiency of full lineups. For example, [4] take a resource-based view for choosing optimal lineups in soccer and describe the type of “human resources” that lead to superior value creation on a team. [5] introduce a skills plus-minus to measure how good individuals are at certain aspects of basketball such as ballhandling, rebounding, and scoring. The authors then compute optimal lineups by measuring synergies and finding players with complimentary skill sets. [6] discuss how APM can be used directly for ranking full lineups. Besides summing over individual player effects in a given lineup, the authors propose a separate parameter for each lineup, constituting a lineup effect. [7] introduce a Bayesian method for quantifying the high degree of uncertainty in individual and lineup rankings.
There are a few existing ranking methods that use techniques from network theory. LinNet was introduced in [8] as an approach to ranking lineups via network embeddings. A network is constructed in which the lineups are nodes, a directed edge $i\rightarrow j$ exists if lineup $i$ outscored lineup $j$, and the corresponding scoring margin is the weight of the edge. For each lineup, the authors compute the average APM of the players in the lineup. The authors then embed this lineup network using node2vec and the resulting latent vectors are used as covariates in logistic regression. The regression response is the difference in average APM for two nodes, which allows the authors to infer the probability that one lineup outperforms another. Similarly, [9] use lineup networks to predict the matchup result between opposing lineups. The authors introduce Inverse Square Metric as a measure for lineup performance prediction, with $\text{ISM}(i,j)~{}=~{}\text{D}(i)\text{D}(j)/|\text{LP}(i,j)|^{2}$, where $\text{D}(i)$ is the degree of lineup $i$ and $\text{LP}(i,j)$ is the length of the shortest path between lineups $i$ and $j$. The authors demonstrate that ISM is more predictive of performance than the node2vec embeddings.
@@
The method most similar to ours is [10], which takes a spectral approach. The authors define a success function $f(L)=\text{pm}_{L}$, where $L$ is one of the $\binom{15}{5}$ 5-person lineups from a roster of 15 and $\text{pm}_{L}$ is the lineup’s plus-minus. The success function is shown to decompose orthogonally into the generalized lineups, hence providing a natural measure of contribution from each generalized lineup. Inference is performed by relating the success function to the Johnson graph. To account for the number of possessions a lineup plays together, the authors normalize their results by the log of the number of possessions played by each lineup, which they call spectral contribution per log possession (SCLP). There is not a natural means to adjusting for opponent with SCLP.
In contrast to SCLP, we utilize a model-based approach via the familiar tools of regression analysis. Our results, as fitted values or coefficient estimates, are easy to interpret and can provide standard error estimates and confidence statements. Furthermore, whereas SCLP requires data from a stable 15-person roster, our methods can be utilized to make predictions for unused lineups and can consider teammates that are added, dropped, or traded throughout the season. While both metrics provide within team rankings, we show that our results nevertheless provide an accurate comparison of generalized lineups across the league. Lastly, as noted by the authors of [10], the various methods of rating players and generalized lineups can be seen as complementary. Each approach offers distinct insights, and when combined, they enhance decision-making by providing a more comprehensive understanding of team and player performance.
@@ 3 Adjusted plus-minus
To illustrate APM, imagine a team with 5 players (labeled A-E), and a 10-minute game with lineups of size 3. The results of one such game are summarized in Table 1. In the linear model described above, the player columns (A:E) define the design matrix $X$ and ‘Score Dif’ is the response vector $Y$. We see that the first lineup (A, B, C) was the most successful with a score differential of +3, indicating that this lineup outscored the other team by 3 total points in the 3 minutes that A, B, and C played together. We also see that Player A has the top ranking PM score of +5, which can be computed via the dot product $X_{\cdot,A}\cdot Y$. However, Player A always played with Player B, so it is unclear which of the two players deserves more credit for their lineups’ successes.
@@
More advanced APM metrics incorporate both home and away players, use a penalized regression setup to stabilize the estimates [11], and incorporate playing time and additional box score statistics [3]. Continuing with our prior example, we can employ a penalized regression scheme that prevents overfitting, and include a diagonal weight matrix of Time, $W$, that takes into account the total time each lineup was active. For instance, ridge regression [12] yields the APM solutions $\widehat{\beta_{r}}=(X^{T}WX+\lambda I)^{-1}X^{T}WY$. In our example, after using cross-validation to choose an appropriate value of the penalization parameter $\lambda$, this results in the improved ranking scheme: A, C, B, D, E.
@@ 3.1 Pairs APM
An intuitive and relatively straightforward approach to account for lower-order combinations in the current construction of APM is to add interaction terms as new columns in the design matrix. Continuing with our toy example, we can add $\binom{5}{2}=10$ additional columns to X, each representing a pair of players. The column vector takes the value 1 when both players are in the line-up and 0 otherwise. See Table 3 for an example. The response vector does not change, and the coefficients obtained from regressing $Y$ on $X$ allows one to simultaneously rank individuals and pairs while accounting for all other players and pairs. We refer to this method as pairs adjusted plus-minus (PAPM).
@@
We note that this idea was alluded to in [2]:
@@ 4 Hypergraphs
In this section, we introduce networks, hypergraphs and line graphs, which are essential to our proposed methods. We begin by connecting the design matrix of APM to a $k$-regular hypergraph. We then introduce a generalization of this to capture all generalized lineups. Finally, we describe our two new methods.
@@ 4.1 Background
Graphs, or networks, are relational data objects composed of nodes or vertices that are linked by edges. Edges represent various types of relationships or interactions between nodes, and a variety of statistical methods can be utilized to model and understand the complex dependencies captured by the network [13]. Given that team lineups are inherently interconnected, networks of game/team data are a natural representation to consider when ranking generalized lineups.
@@ 4.2.1 Hypergraph APM
Using HAPM, can we rank combinations of players? The answer is yes. In addition to the regression coefficients, which correspond to the players/columns of our design matrix, we have the predicted values $\widehat{y}_{\text{HAPM}}$ that correspond to the combinations/rows. Therefore, we have a natural solution to rank combinations of any size. For example, if we want to find the best combination of $m$ players, we sort the predicted values $\widehat{y}_{\text{HAPM}}$ among all combinations whose row sums of $X_{\text{HAPM}}$ are equal to $m$. We can use this same approach to separately rank individuals, pairs, etc.
@@
For individuals, we can use either the coefficients $\widehat{\beta}_{\text{HAPM}}$ or the predicted values $\widehat{y}_{\text{HAPM}}$ corresponding to row sums of $X_{\text{HAPM}}$ equal to 1. We show in the results that these are highly correlated, so there is minimal difference in using one or the other.
@@ 4.2.2 Line graph APM
HAPM utilizes the hypergraph representation implicitly, yet it does not explicitly consider the connectivity patterns within the hypergraph. To address this, we draw inspiration from functional data analysis [14] and propose a Bayesian regression approach. This approach resembles kernel regression on a graph [15], allowing us to extend regression-based techniques to graph-indexed data, where conventional models and assumptions may not be applicable. In our case, it allows for a ranking of individual players and generalized lineups while considering all active combinations of players on a team.
@@
showing that the term penalizes the weighted sum of squared differences for coefficients of adjacent vertices in the network [13]. Moreover, pairs of nodes with larger edge weights, or more similar generalized lineups, will be forced to have more similar predicted values. Given that $L_{w}$ is a symmetric matrix, it realizes the eigendecomposition $L_{w}=\Phi\Xi\Phi$, and we can therefore represent $\beta$ using a basis expansion with respect to the eigenvectors of $L_{w}$ by letting $\beta=\Phi\theta$. In short, we use the eigenvectors of the Laplacian as predictor variables, find a set of corresponding coefficients $\hat{\theta}$ through Bayesian linear regression, and then perform matrix-vector multiplication to solve for the vertex-indexed intercepts $\widehat{\beta_{v}}=\widehat{Y_{v}}.$
We do not have to use the full set of eigenvectors to represent $\beta$. Instead, we can extract most of the information in $L_{w}$ through a low-order representation such that $\beta=\Phi_{1:\tau}\theta$, where $\Phi_{1:\tau}$ represents the first $\tau$ eigenvectors (ordered by the eigenvalues $\xi_{1}<\dots<\xi_{\tau}$). In practice, we select $\tau$ based on an ‘elbow’ method, examining sequential differences in the size of ${\xi_{i}}$. Lastly, we select $\lambda$ using a grid search, selecting a final value of $\lambda$ that produces fitted values that range over comparable values to that of the observed data.
@@ 5.2 Rank Correlation
Evaluating player rankings is ultimately qualitative since there is no notion of ground truth. However, as a baseline comparison, we can see how closely our rankings are correlated with those based on advanced metrics such as box score plus-minus (BPM), value over replacement player (VORP), player efficiency rating (PER), and win shares per 48 minutes (WS.48) [16]. Ideally, the individual rankings we obtain from our methods, while also providing rankings for all combinations of players, will be in general agreement with known methods designed to rank individuals.
@@
For each of the advanced metrics, we see that HAPM is better (i.e. in more agreement with the advanced metrics) than APM. Similarly, LAPM is as good or better than APM with respect to BPM, VORP, and WS.48; the only metric that APM outperforms LAPM is PER, which is known to be mostly a measure of offensive production or scoring efficiency and therefore a less well-rounded metric. LAPM is most highly correlated with WS.48. This may be because WS.48 considers offensive and defensive contributions, incorporates playing time, and is highly connected to the corresponding team’s overall performance and win record. As discussed below, HAPM and LAPM, when aggregated across the league, are both highly correlated with win percentages.
In Figure 3, we can also see what is lost when moving away from traditional APM, which is modeled at the league level and adjusts for defense. Moving left to right among the methods, we see that PAPM is generally worse than APM for ranking individuals, which is not surprising since including additional pair columns in the design matrix requires greater regularization (i.e. shrinkage of all of the coefficients). We also see that computing APM at the team level is a slight decrease compared to the league level, and computing PAPM at the team level leads to a considerable drop in performance.
From these comparisons we see that both HAPM and LAPM align more closely with advanced metrics compared to APM, despite neither explicitly incorporating defensive considerations nor being computed at the league level. Moreover, HAPM and LAPM offer simultaneous rankings for all generalized lineups.

### v2

@@ 1 Introduction
One of the most common approaches to measuring an individual’s effect in team sports is box-score plus-minus (PM). Originally used in hockey, a player’s PM is their team’s cumulative difference in score when that player is in the game. While intuitively appealing, PM is sensitive to substitution patterns. For instance, one very good player can artificially inflate the PM of their teammates, and quantifying this effect becomes challenging when certain combinations of players consistently share playing time. The most common approach to account for this is known as adjusted PM (APM). APM was developed for basketball (Rosenbaum, 2004) and hockey (Macdonald, 2011) as a method for comparing individual players while accounting for those that play together. We provide the mathematical details of APM in Section 3.
The intention of APM and its variants (Hvattum, 2019) is to isolate an individual player’s effect. However, this approach fails to capture any interaction effects between players. Experience, complementary skill sets, and general synergy often result in a group of players whose combined effect surpasses the sum of their individual contributions. This phenomena is in part why substitution patterns are not uniformly random, and why managers and coaches play a crucial role in the success of a team. Likewise, when assessing and ranking an individual’s contribution to a team’s success, it is critical to account for the player’s ability to create synergies with the various combinations of their teammates. We refer to these various combinations (individuals, pairs, trios, etc.) as generalized lineups and propose a novel approach to quantify the contribution of all generalized lineups for a team simultaneously.
@@
Both of these approaches have obvious drawbacks, and neither is able to rank the generalized lineups while accounting for interaction effects. To fill the gap between these two extreme solutions, we introduce a new class of methods for ranking lineups and players that adjusts not only for a full lineup, but all lower-order combinations of players. Our methods are motivated by the hypergraph representation of a lineup, which we show is a generalization of the design matrix used in the ordinary APM method. The result is a unified approach for simultaneously ranking any generalized lineup including individuals, pairs, and full lineups. Furthermore, our methods lend themselves to interpretable visualizations that allow one to quickly identify positive and negative synergies among groups of players on a team.
@@ 2 Related Work
While methods for ranking lower-order combinations of players and full lineups simultaneously are lacking, there has been a great deal of research on measuring the efficiency of full lineups. For example, Lechner and Vidar Gudmundsson (2012) take a resource-based view for choosing optimal lineups in soccer and describe the type of “human resources” that lead to superior value creation on a team. Maymin et al. (2013) introduce a skills plus-minus to measure how good individuals are at certain aspects of basketball such as ballhandling, rebounding, and scoring. The authors then compute optimal lineups by measuring synergies and finding players with complementary skill sets. Grassetti et al. (2021) discuss how APM can be used directly for ranking full lineups: Besides summing over individual player effects in a given lineup, the authors propose a separate parameter for each lineup, constituting a lineup effect. Barrientos et al. (2023) introduce a Bayesian method for quantifying the high degree of uncertainty in individual and lineup rankings.
There are a few existing ranking methods that use techniques from network theory. LinNet was introduced in Pelechrinis (2018) as an approach to ranking lineups via network embeddings. A network is constructed in which the lineups are nodes, a directed edge $i\rightarrow j$ exists if lineup $i$ outscored lineup $j$, and the corresponding scoring margin is the weight of the edge. For each lineup, the authors compute the average APM of the players in the lineup. The authors then embed this lineup network using node2vec and the resulting latent vectors are used as covariates in logistic regression. The regression response is the difference in average APM for two nodes, which allows the authors to infer the probability that one lineup outperforms another. Similarly, Ahmadalinezhad and Makrehchi (2020) use lineup networks to predict the matchup result between opposing lineups. The authors introduce Inverse Square Metric as a measure for lineup performance prediction, with $\text{ISM}(i,j)~{}=~{}\text{D}(i)\text{D}(j)/|\text{LP}(i,j)|^{2}$, where $\text{D}(i)$ is the degree of lineup $i$ and $\text{LP}(i,j)$ is the length of the shortest path between lineups $i$ and $j$. The authors demonstrate that ISM is more predictive of performance than the node2vec embeddings.
@@
The method most similar to ours is Devlin and Uminsky (2020), which takes a spectral approach. The authors define a success function $f(L)=\text{pm}_{L}$, where $L$ is one of the $\binom{15}{5}$ 5-person lineups from a roster of 15 and $\text{pm}_{L}$ is the lineup’s plus-minus. The success function is shown to decompose orthogonally into the generalized lineups, hence providing a natural measure of contribution from each generalized lineup. Inference is performed by relating the success function to the Johnson graph. To account for the number of possessions a lineup plays together, the authors normalize their results by the log of the number of possessions played by each lineup, which they call spectral contribution per log possession (SCLP). There is not a natural means to adjusting for opponent with SCLP.
In contrast to SCLP, we utilize a model-based approach via the familiar tools of regression analysis. Our results, as fitted values or coefficient estimates, are easy to interpret and can provide standard error estimates and confidence statements. Furthermore, whereas SCLP requires data from a stable 15-person roster, our methods can be utilized to make predictions for unused lineups and can consider teammates that are added, dropped, or traded throughout the season. While both metrics provide within team rankings, we show that our results nevertheless provide an accurate comparison of generalized lineups across the league. Lastly, as noted by the authors of Devlin and Uminsky (2020), the various methods of rating players and generalized lineups can be seen as complementary: Each approach offers distinct insights, and when combined, they enhance decision-making by providing a more comprehensive understanding of team and player performance.
@@ 3 Adjusted plus-minus
To illustrate APM, imagine a team with 5 players labeled A-E, and a 10-minute game with lineups of size 3. The results of one such game are summarized in Table 1. In the linear model described above, the player columns (A:E) define the design matrix $X$ and ‘Score Dif’ is the response vector $Y$. We see that the first lineup (A, B, C) was the most successful with a score differential of +3, indicating that this lineup outscored the other team by 3 total points in the 3 minutes that A, B, and C played together. We also see that Player A has the top ranking PM score of +5, which can be computed via the dot product $X_{\cdot,A}\cdot Y$. However, Player A always played with Player B, so it is unclear which of the two players deserves more credit for their lineups’ successes.
@@
More advanced APM metrics incorporate both home and away players, use a penalized regression setup to stabilize the estimates (Sill, 2010), and incorporate playing time and additional box score statistics (Hvattum, 2019). Continuing with our prior example, we can employ a penalized regression scheme that prevents overfitting, and include a diagonal weight matrix of Time, $W$, that takes into account the total time each lineup was active. For instance, ridge regression (Hoerl and Kennard, 1970) yields the APM solutions $\widehat{\beta_{r}}=(X^{T}WX+\lambda I)^{-1}X^{T}WY$. In our example, after using cross-validation to choose an appropriate value of the penalization parameter $\lambda$, this results in the improved ranking scheme: A, C, B, D, E. Note that each $\widehat{\beta}$ is an estimated parameter, so we could quantify the uncertainty and provide the implied rank intervals. We omit this here, for clarity, but discuss it later in Section 5 with our application to real data.
@@ 3.1 Pairs APM
An intuitive and relatively straightforward approach to account for lower-order combinations in the current construction of APM is to add interaction terms as new columns in the design matrix. Continuing with our toy example, we can add $\binom{5}{2}=10$ additional columns to X, each representing a pair of players. The column vector takes the value 1 when both players are in the line-up and 0 otherwise. See Table 3 for an example. The response vector does not change, and the coefficients obtained from regressing $Y$ on $X$ allow one to simultaneously rank individuals and pairs while accounting for all other players and pairs. We refer to this method as pairs adjusted plus-minus (PAPM).
@@
We note that this idea was alluded to in Macdonald (2011):
@@ 4 Hypergraphs
In this section, we introduce networks, hypergraphs, and line graphs, which are essential to our proposed methods. We begin by connecting the design matrix of APM to a $k$-regular hypergraph. We then introduce a generalization of this to capture all generalized lineups. Finally, we describe our two new methods.
@@ 4.1 Background
Graphs, or networks, are relational data objects composed of nodes or vertices that are linked by edges. Edges represent various types of relationships or interactions between nodes, and a variety of statistical methods can be utilized to model and understand the complex dependencies captured by the network (Kolaczyk and Csárdi, 2014). Given that team lineups are inherently interconnected, networks of game/team data are a natural representation to consider when ranking generalized lineups.
@@ 4.2.1 Hypergraph APM
Using HAPM, can we rank combinations of players? The answer is yes. In addition to the regression coefficients, which correspond to the players/columns of our design matrix, we have the predicted values $\widehat{y}_{\text{HAPM}}$ that correspond to the combinations/rows. Therefore, we have a natural solution to rank combinations of any size. For example, if we want to find the best combination of $m$ players, we sort the predicted values $\widehat{y}_{\text{HAPM}}$ among all combinations whose row sums of $X_{\text{HAPM}}$ are equal to $m$. We can use this same approach to separately rank pairs, trios, etc.
@@
For individuals, we can use either the coefficients $\widehat{\beta}$ or the predicted values $\widehat{y}$ corresponding to row sums of $X$ equal to 1. These values are equivalent because $\widehat{y}=X\widehat{\beta}$ and for an individual $i$, there is only one nonzero column, $X_{\cdot,i}$, corresponding to that individual, hence $\widehat{y}_{i}=X_{\cdot,i}\widehat{\beta}=\widehat{\beta}_{i}$.
@@ 4.2.2 Line graph APM
HAPM utilizes the hypergraph representation implicitly, yet it does not explicitly consider the connectivity patterns within the hypergraph. To address this, we draw inspiration from functional data analysis (Ramsay and Silverman, 2002) and propose a Bayesian regression approach. This approach resembles kernel regression on a graph (Smola and Kondor, 2003), allowing us to extend regression-based techniques to graph-indexed data, where conventional models and assumptions may not be applicable. In our case, it allows for a ranking of individual players and generalized lineups while considering all active combinations of players on a team.
@@
showing that the term penalizes the weighted sum of squared differences for coefficients of adjacent vertices in the network (Kolaczyk and Csárdi, 2014). Moreover, pairs of nodes with larger edge weights, or more similar generalized lineups, will be forced to have more similar predicted values. Given that $L_{w}$ is a symmetric matrix, it realizes the eigendecomposition $L_{w}=\Phi\Xi\Phi$, and we can therefore represent $\beta$ using a basis expansion with respect to the eigenvectors of $L_{w}$ by letting $\beta=\Phi\theta$ (Upton and Carvalho, 2017). In short, we use the eigenvectors of the Laplacian as predictor variables, find a set of corresponding coefficients $\hat{\theta}$ through Bayesian linear regression, and then perform matrix-vector multiplication to solve for the vertex-indexed intercepts $\widehat{\beta_{v}}=\widehat{Y_{v}}.$
We do not have to use the full set of eigenvectors to represent $\beta$. Instead, we can extract most of the information in $L_{w}$ through a low-order representation such that $\beta=\Phi_{1:\tau}\theta$, where $\Phi_{1:\tau}$ represents the first $\tau$ eigenvectors (ordered by the eigenvalues $\xi_{1}<\dots<\xi_{\tau}$). In practice, we select $\tau$ based on an ‘elbow’ method, examining sequential differences in the size of ${\xi_{i}}$.
@@ 5.1 Data Processing
Finally, for the case study analysis, we explore the variability in the fitted values of HAPM and LAPM, along with the implied uncertainty in their rankings. For HAPM, we bootstrap by sampling with replacement the rows of the design matrix. For each bootstrap sample, we fit a ridge regression model and use the model to make predictions on all generalized lineups. We examine and present the variation in the fitted HAPM values across the bootstrapped samples. We also rank the fitted values within combinations of the same size (individuals, pairs, etc.), providing an estimate of rank uncertainty. For LAPM, we sample from the posterior distributions of the coefficients utilizing a Metropolis-adjusted Langevin algorithm. The details are provided in Appendix A.
@@ 5.2 Rank Correlation
Evaluating player rankings is ultimately qualitative since there is no notion of ground truth. However, as a baseline comparison, we can see how closely our rankings are correlated with those based on advanced metrics such as box score plus-minus (BPM), value over replacement player (VORP), player efficiency rating (PER), and win shares per 48 minutes (WS.48) (Basketball-Reference, 2024). Ideally, the individual rankings we obtain from our methods will be in general agreement with known methods designed to rank individuals, while also providing rankings for all combinations of players.
@@
For all of the ridge regression models (APM, PAPM, and HAPM), we chose the regularization parameter $\lambda$ that “gives the most regularized model such that the cross-validated error is within one standard error of the minimum,” which is the default choice for the glmnet package (Hastie et al., 2021). We explored other criterion for choosing $\lambda$, but this valued maximized the correlation for APM and PAPM. On the other hand, HAPM was robust to different choices of $\lambda$.
For each of the advanced metrics, other than WS.48, we see that HAPM is better (i.e. in more agreement with the advanced metrics) than APM. In contrast, LAPM is most highly correlated with WS.48. This is important because WS.48 considers offensive and defensive contributions (rather than PER, which is primarily a measure of offensive production and scoring efficiency), incorporates playing time, and is highly connected to the corresponding team’s overall performance and win record.
From these comparisons, we see that HAPM and LAPM align with advanced metrics better than APM despite neither explicitly incorporating defensive considerations nor being computed at the league level. Moreover, HAPM and LAPM offer simultaneous rankings for all generalized lineups and are more robust to choice of penalization parameter.